# Quantum cohomology of projective bundles

### v3

@@ 2.3. The vector bundle case
with $\alpha_{1},\dots,\alpha_{n}\in H^{*}_{\mathbb{C}^{\times}}(V)=H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ by the virtual localization formula [22], where $\lambda$ is the $\mathbb{C}^{\times}$-equivariant (or $S^{1}$-equivariant) parameter [^1] In this paper, all $S^{1}$-actions are induced from $\mathbb{C}^{\times}$-actions, so we use the terms ‘$S^{1}$-equivariant’ and ‘$\mathbb{C}^{\times}$-equivariant’ interchangeably.. The equivariant Gromov-Witten invariants take values in $\mathbb{C}[\lambda,\lambda^{-1}]$. In §2.5 below, we describe them as the Gromov-Witten invariants of $B$ twisted by $V$ and the inverse equivariant Euler class.
@@
where $K[\![\boldsymbol{\tau}]\!]$ denotes the space of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}$ with coefficients in $K$ for a module $K$ (we use the boldface symbol $\boldsymbol{\tau}$ to distinguish infinitely many parameters $\{\tau^{i,k}\}$ from finitely many ones $\{\tau^{i}\}$).
@@ 2.4. Givental Lagrangian cones
where $\mathbf{t}(z)=\sum_{n\geq 0}t_{n}z^{n}$ with $t_{n}=\sum_{i=0}^{s}t^{i}_{n}\phi_{i}\in H^{*}(X)$. The argument $-z+\mathbf{t}(z)$ represents a point in the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$ and $\mathcal{F}_{X}$ is regarded as a function on the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$. The Givental Lagrangian cone $\mathcal{L}_{X}^{\rm orig}\subset\mathcal{H}_{X}$ is a submanifold-germ defined as the graph of $d\mathcal{F}_{X}$ in $T^{*}\mathcal{H}_{+}\cong\mathcal{H}_{X}$. It consists of
(2.14) $-z+\mathbf{t}(z)+\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\frac{\partial\mathcal{F}_{X}}{\partial t_{k}^{i}}\\ =-z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{X}\frac{Q^{d}}{n!}$
@@ 2.5. Twisted Gromov-Witten invariants and quantum Riemann-Roch theorem
    $W$ is convex, i.e. we have $H^{1}(C,f^{*}W)=0$ for any genus-zero stable maps $C\to X$ and $\mathbf{c}$ is the equivariant Euler class $e_{\lambda}$. This corresponds to the choice of parameters
@@
where $i\colon Z\to X$ is the inclusion map. Moreover, the fundamental solution $M_{X,(W,e_{\lambda})}$ in the $(W,e_{\lambda})$-twisted theory satisfies (see [44, §2.1], [26, Proposition 2.4])
@@ 2.6. Moving points on the Givental cone via differential operators
This can be shown easily by induction on powers of $Q$ and $y$ (see [23, Theorem 3.17] for a related result). Then we have
@@
We construct $w_{i}$ by induction on powers of $Q$ and $y$. Take an ample class $\omega$ of $X$, and consider the additive valuation $v\colon\mathbb{C}[\![Q,y]\!]\to\mathbb{R}\cup\{\infty\}$ given by $v(Q^{d})=\omega\cdot d$ and $v(y_{i})=i$. Suppose by induction that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that the equation (2.20) holds modulo terms with valuation $>v_{0}$ for some $v_{0}\in v(\mathbb{C}[\![Q,y]\!])$. Let $v_{1}>v_{0}$ be the next possible value of the valuation $v$. Write $\exp(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}})g^{\prime}=1-\sum_{i=0}^{s}\delta_{i}\phi_{i}$ for some $\delta_{i}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ with $v(\delta_{i})\geq v_{1}$. Using the fact that $z\nabla_{\tau^{i}}1=\phi_{i}$, we have that $\exp\left(\sum_{i=0}^{s}(w_{i}+\delta_{i})z\nabla_{\tau^{i}}\right)g^{\prime}$ equals 1 modulo terms with valuation $>v_{1}$. This completes the induction and the proof of the claim. Setting $G=\sum_{i=0}^{s}w_{i}z\partial_{\tau^{i}}$, we have
@@ 2.7. Shift operators
In this section, we work with the $\mathbb{C}$-basis $\{\phi_{i}\lambda^{k}\}$ and the associated $\mathbb{C}$-linear coordinates $\{\tau^{i,k}\}$ of $H^{*}_{S^{1}}(V)$ as in §2.3. The shift operator shall be defined over the ring $\mathbb{C}[\![\boldsymbol{\tau}]\!]$ of formal power series in $\{\tau^{i,k}\}$. Since $V$ is semiprojective and the fibrewise scalar $S^{1}$-action is seminegative [^5] On the other hand, the inverse $S^{1}$-action is not seminegative: this is the reason why $\mathbb{S}(\tau)^{-1}$ is not defined without localization. (in the sense of [27, Definition 3.3]), the $S^{1}$-action defines the shift operator [27, Definition 3.9]
@@
$[\partial_{\tau^{i,k}},\mathcal{S}]=[\xi Q\partial_{Q}+z^{-1}\xi,\mathcal{S}]=[z\partial_{z}+z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V},\mathcal{S}]=0.$
@@
The image of $\mathbb{S}(\tau)$ is a sub $D$-module since $\mathbb{S}(\tau)$ commutes with the connection. This fact is related to the quantum Serre [7, Corollary 9]: the above submodule can be identified with the quantum $D$-module of the $(V^{\vee},e_{\lambda})$-twisted theory [30, Theorems 2.11, 3.14].
@@ 3. Proof of the mirror theorem
In this section we prove Theorem 1.1 (and a more general version Theorem 3.3 below). Let $\mathbb{P}(V)\to B$ be the projective bundle associated with a vector bundle $V\to B$ of rank $r\geq 2$. Let $p=c_{1}(\mathcal{O}(1))$ be the first Chern class of the relative $\mathcal{O}(1)$ over $\mathbb{P}(V)$. This class defines the following splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})$:
@@
We write $Q$ for the Novikov variable for $B$, and $q$ for the Novikov variable for $\mathbb{P}(V)$ associated with the class of a line in a fibre. Via the above splitting (3.1), we identify $(q,Q)$ with the Novikov variables for $\mathbb{P}(V)$: a curve class $(k,d)\in\mathbb{Z}\oplus H_{2}(B,\mathbb{Z})\cong H_{2}(\mathbb{P}(V),\mathbb{Z})$ corresponds to the monomial $q^{k}Q^{d}$.
@@ 3.1. Outline
(3.2) $0\to V\to\mathcal{O}^{\oplus N}\to\mathcal{Q}\to 0$
@@
Our starting point is the fact that Theorem 3.3 holds for the trivial bundle $B\times\mathbb{C}^{N}$: the $J$-function of $B\times\mathbb{P}^{N-1}$ at parameter $\tau+tp$ (with $\tau\in H^{*}(B)$) is given by
$J_{B\times\mathbb{P}^{N-1}}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}J_{B}(\tau).$
@@
3.2. Informal proof
We first give a slightly informal proof, ignoring the issue on adic convergence. A rigorous proof follows in the next section, but it is only technically complicated and we recommend the reader to begin with the discussion in this section. We introduce a modified QRR operator:
(3.3) $\Delta^{\lambda}_{W}:=e^{\operatorname{rank}(W)(\lambda\log\lambda-\lambda)/z}\Delta_{(W,e_{\lambda}^{-1})}$
where $\Delta_{(W,e_{\lambda}^{-1})}$ is the QRR operator (2.19) of the inverse Euler twist. This arises from the asymptotic expansion (Stirling’s approximation) of the $\Gamma$-function:
$\Delta^{\lambda}_{W}\sim\prod_{\delta}\sqrt{\frac{z}{2\pi}}z^{\frac{\lambda+\delta}{z}}\Gamma\left(1+\frac{\lambda+\delta}{z}\right)$
where $\delta$ ranges over the Chern roots of $W$. By the String Equation, the Givental cone is preserved by the action of $e^{g(z)/z}$, with $g(z)$ a nonnegative power series of $z$ with scalar coefficients. Hence we can use $\Delta_{W}^{\lambda}$ instead of $\Delta_{(W,e_{\lambda}^{-1})}$ in the QRR Theorem 2.5. The identity $\Gamma(1+x)=x\Gamma(x)$ implies that
@@ 3.2. Informal proof
The exact sequence (3.2) implies $\Delta_{V}^{\lambda}\Delta_{\mathcal{Q}}^{\lambda}=\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$. Since $\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$ preserves the Givental cone (again by the String Equation), we have $\Delta_{\mathcal{Q}}^{\lambda}\mathcal{L}_{B,(V,e_{\lambda}^{-1})}=\mathcal{L}_{B}$ by the QRR Theorem 2.5. Let $J_{V}^{\lambda}(\tau)$ be the equivariant $J$-function of $V$ so that $zJ_{V}^{\lambda}(\tau)\in\mathcal{L}_{B,(V,e_{\lambda}^{-1})}$. Then $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)$ lies in $\mathcal{L}_{B}$. As in Lemma 2.8, we expect to have a differential operator $F(\lambda)=F(\lambda;\tau,z\partial_{\tau},z)$ in $\tau$ such that
(3.5) $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)=e^{F(\lambda)/z}zJ_{B}(\tau).$
This is not justified by the argument in the proof of Lemma 2.8 due to the lack of an appropriate formal topology; $\Delta^{\lambda}_{\mathcal{Q}}$ contains both positive and negative powers of $\lambda$ and also $\log\lambda$. Let us proceed assuming that we have such an $F(\lambda)$.
Consider the differential operator $\exp(F(\lambda+z\partial_{t})/z)$ in $t$ and $\tau$; here we expand $F(\lambda+z\partial_{t})$ in Taylor series of $z\partial_{t}$. By Lemma 2.7, $e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$ lies in $\mathcal{L}_{B\times\mathbb{P}^{N-1}}$. Therefore
$I^{\lambda}(\tau,t):=(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$
lies in the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$ by the QRR Theorem 2.5. We calculate
$I^{\lambda}(\tau,t)$ $=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+p+kz)/z}J_{B}(\tau)$
$=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}}^{\lambda+p})^{-1}\Delta^{\lambda+p+kz}_{\mathcal{Q}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.5)
@@
(3.6) $=\sum_{k\geq 0}e^{tp/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+p+\epsilon+cz)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.4)
where $\epsilon$ ranges over the Chern roots of $\mathcal{Q}$ in the last line. As discussed in §2.5, $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ lies in the Givental cone $\mathcal{L}_{\mathbb{P}(V)}$ of $\mathbb{P}(V)$, where $i\colon\mathbb{P}(V)\hookrightarrow B\times\mathbb{P}^{N-1}$ is the inclusion. This gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1. Note that $i_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\to H_{2}(B\times\mathbb{P}^{N-1},\mathbb{Z})$ is an isomorphism whenever $r=\operatorname{rank}(V)\geq 2$.
@@
Another point that we glossed over in the above discussion is that the modified Riemann-Roch operator (3.3)
(3.7) $\Delta^{\lambda}_{W}=\prod_{\delta}e^{\frac{(\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta)}{z}+\frac{1}{2}\log(\lambda+\delta)+\sum_{n\geq 2}\frac{B_{n}}{n(n-1)}\left(\frac{z}{\lambda+\delta}\right)^{n-1}}$

### v4

@@ 2.3. The vector bundle case
with $\alpha_{1},\dots,\alpha_{n}\in H^{*}_{\mathbb{C}^{\times}}(V)=H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ by the virtual localization formula [22], where $\lambda$ is the $\mathbb{C}^{\times}$-equivariant (or $S^{1}$-equivariant) parameter [^3] In this paper, all $S^{1}$-actions are induced from $\mathbb{C}^{\times}$-actions, so we use the terms $S^{1}$-equivariant’ and $\mathbb{C}^{\times}$-equivariant’ interchangeably.. The equivariant Gromov-Witten invariants take values in $\mathbb{C}[\lambda,\lambda^{-1}]$. In §2.5 below, we describe them as the Gromov-Witten invariants of $B$ twisted by $V$ and the inverse equivariant Euler class.
@@
where $K[\![\boldsymbol{\tau}]\!]$ denotes the space of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$ with coefficients in $K$ for a module $K$ (we use the boldface symbol $\boldsymbol{\tau}$ to distinguish infinitely many parameters $\{\tau^{i,k}\}$ from finitely many ones $\{\tau^{i}\}$).
@@ 2.4. Givental Lagrangian cones
where $\mathbf{t}(z)=\sum_{n\geq 0}t_{n}z^{n}$ with $t_{n}=\sum_{i=0}^{s}t^{i}_{n}\phi_{i}\in H^{*}(X)$. The argument $z+\mathbf{t}(z)$ represents a point in the formal neighbourhood of z-z in $\mathcal{H}_{+}$ and $\mathcal{F}_{X}$ is regarded as a function on the formal neighbourhood of z-z in $\mathcal{H}_{+}$. The Givental Lagrangian cone $\mathcal{L}_{X}^{\rm orig}\subset\mathcal{H}_{X}$ is a submanifold-germ defined as the graph of $d\mathcal{F}_{X}$ in $T^{*}\mathcal{H}_{+}\cong\mathcal{H}_{X}$. It consists of
(2.14) $z+\mathbf{t}(z)+\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\frac{\partial\mathcal{F}_{X}}{\partial t_{k}^{i}}\\ =-z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{X}\frac{Q^{d}}{n!}$
@@ 2.5. Twisted Gromov-Witten invariants and quantum Riemann-Roch theorem
    $W$ is convex, i.e. we have $H^{1}(C,f^{*}W)=0$ for any genus-zero stable map $C\to X$ and $\mathbf{c}$ is the equivariant Euler class $e_{\lambda}$. This corresponds to the choice of parameters
@@
where $i\colon Z\to X$ is the inclusion map. Moreover, the fundamental solution $M_{X,(W,e_{\lambda})}$ in the $(W,e_{\lambda})$-twisted theory satisfies (see [47, §2.1], [26, Proposition 2.4])
@@ 2.6. Moving points on the Givental cone via differential operators
This can be shown easily by induction on powers of $Q$ and $y$ (see [24, Theorem 3.17] for a related result). Then we have
@@
We construct $w_{i}$ by induction on powers of $Q$ and $y$. Take an ample class $\omega$ of $X$, and consider the additive valuation $v\colon\mathbb{C}[\![Q,y]\!]\to\mathbb{R}\cup\{\infty\}$ given by $v(Q^{d})=\omega\cdot d$ and $v(y_{i})=i$. Suppose by induction that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that the equation (2.20) holds modulo terms with valuation $v_{0}$ for some $v_{0}\in v(\mathbb{C}[\![Q,y]\!])$. Let $v_{1}>v_{0}$ be the next possible value of the valuation $v$. Write $\exp(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}})g^{\prime}=1-\sum_{i=0}^{s}\delta_{i}\phi_{i}$ for some $\delta_{i}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ with $v(\delta_{i})\geq v_{1}$. Using the fact that $z\nabla_{\tau^{i}}1=\phi_{i}$, we have that $\exp\left(\sum_{i=0}^{s}(w_{i}+\delta_{i})z\nabla_{\tau^{i}}\right)g^{\prime}$ equals 1 modulo terms with valuation $v_{1}$. This completes the induction and the proof of the claim. Setting $G=\sum_{i=0}^{s}w_{i}z\partial_{\tau^{i}}$, we have
@@ 2.7. Shift operators
In this section, we work with the $\mathbb{C}$-basis $\{\phi_{i}\lambda^{k}\}$ and the associated $\mathbb{C}$-linear coordinates $\{\tau^{i,k}\}$ of $H^{*}_{S^{1}}(V)$ as in §2.3. The shift operator shall be defined over the ring $\mathbb{C}[\![\boldsymbol{\tau}]\!]$ of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$. Since $V$ is semi-projective and the fibrewise scalar $S^{1}$-action is seminegative [^7] On the other hand, the inverse $S^{1}$-action is not seminegative: this is the reason why $\mathbb{S}(\tau)^{-1}$ is not defined without localization. (in the sense of [27, Definition 3.3]), the $S^{1}$-action defines the shift operator [27, Definition 3.9]
@@
$\partial_{\tau^{i,k}},\mathcal{S}]=[\xi Q\partial_{Q}+z^{-1}\xi,\mathcal{S}]=[z\partial_{z}+z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V},\mathcal{S}]=0.$
@@
The image of $\mathbb{S}(\tau)$ is a sub $D$-module since $\mathbb{S}(\tau)$ commutes with the connection. This fact is related to the quantum Serre [7, Corollary 9]: the above submodule can be identified with the quantum $D$-module of the $(V^{\vee},e_{\lambda})$-twisted theory [31, Theorems 2.11, 3.14].
@@ 3. Proof of the mirror theorem
In this section we prove Theorem 1.1 (and a more general version Theorem 3.4 below). Let $\pi\colon\mathbb{P}(V)\to B$ be the projective bundle associated with a vector bundle $V\to B$ of rank $r\geq 2$ such that $V^{\vee}$ is globally generated. Let $p=c_{1}(\mathcal{O}(1))$ be the first Chern class of the relative $\mathcal{O}(1)$ over $\mathbb{P}(V)$. This class defines a splitting of the natural sequence $0\to\mathbb{Z}\to H_{2}(\mathbb{P}(V),\mathbb{Z})\xrightarrow{\pi_{*}}H_{2}(B,\mathbb{Z})\to 0$ as follows:
@@
We write $Q$ for the Novikov variable of the base $B$, and $q$ for the Novikov variable of $\mathbb{P}(V)$ associated with the class of a line in a fibre. Via the above splitting (3.1), we identify the pair $(q,Q)$ with the Novikov variables for $\mathbb{P}(V)$; more precisely, a curve class $d\in H_{2}(\mathbb{P}(V),\mathbb{Z})$ corresponds to the monomial $q^{p\cdot d}Q^{\pi_{*}d}$. The assumption that $V^{\vee}$ is globally generated implies that the class $p$ is nef; consequently the exponent of $q$ in this monomial is nonnegative whenever $d$ is an effective class.
@@ 3.1. Outline
(3.2) $0\to V\to\mathcal{O}^{\oplus N}\to\mathcal{Q}\to 0.$
@@
Our starting point is the fact that Theorem 3.4 holds for the trivial bundle $B\times\mathbb{C}^{N}\to B$: the $J$-function of $B\times\mathbb{P}^{N-1}$ at parameter $\tau+tp$ (with $\tau\in H^{*}(B)$) is given by
(3.3) $J_{B\times\mathbb{P}^{N-1}}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}J_{B}(\tau).$
@@
3.2. Proof of Theorem 1.1
We provide a rigorous treatment based on the $\lambda^{-1}$-adic topology, following the method of Remark 2.6. Throughout this section, $\tau$ is assumed to be a non-equivariant class in $H^{*}(V)=H^{*}(B)$ (i.e. independent of $\lambda$). A generalization to equivariant classes will be discussed in Theorem 3.4.
We introduce a modified Quantum Riemann-Roch (QRR) operator:
(3.4) $\begin{split}\Delta^{\lambda}_{W}&:=e^{\operatorname{rank}(W)(\lambda\log\lambda-\lambda)/z}\Delta_{(W,e_{\lambda}^{-1})}\\ &=\prod_{\delta}e^{\frac{(\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta)}{z}+\frac{1}{2}\log(\lambda+\delta)+\sum_{n\geq 2}\frac{B_{n}}{n(n-1)}\left(\frac{z}{\lambda+\delta}\right)^{n-1}}\end{split}$
where $\Delta_{(W,e_{\lambda}^{-1})}$ is the QRR operator (2.19) of the inverse Euler twist, and $\delta$ ranges over the Chern roots of $W$. This arises from the asymptotic expansion (Stirling’s approximation) of the $\Gamma$-function:
$\Delta^{\lambda}_{W}\sim\prod_{\delta}\sqrt{\frac{z}{2\pi}}z^{\frac{\lambda+\delta}{z}}\Gamma\left(1+\frac{\lambda+\delta}{z}\right).$
The identity $\Gamma(1+x)=x\Gamma(x)$ implies that
@@
Ignoring issues of adic convergence (see Remark 3.3), we may regard the modified operator $\Delta_{W}^{\lambda}$ as playing the same role as the original $\Delta_{(W,e_{\lambda}^{-1})}$, because $e^{c/z}$ preserves the Givental cone for any constant $c$.
Let $\tilde{e}_{\lambda}(\cdot)=\sum_{i\geq 0}\lambda^{-i}c_{i}(\cdot)$ be the variant of $e_{\lambda}$ introduced in Remark 2.6 and let $\widetilde{\Delta}_{W}^{\lambda}:=\Delta_{(W,\tilde{e}_{\lambda}^{-1})}$ denote the QRR operator (2.19) associated with the $(W,\tilde{e}_{\lambda}^{-1})$-twist. We have $\widetilde{\Delta}_{W}^{\lambda}=1+O(\lambda^{-1})$. We also set $\widetilde{J}_{V}^{\lambda}(\tau):=J_{V}^{\lambda}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$; this does not contain positive powers of $\lambda$ and gives the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory by Remark 2.6. We have
$\Delta_{W}^{\lambda}=e^{(\operatorname{rank}(W)\theta(\lambda)+c_{1}(W)\log\lambda)/z}\widetilde{\Delta}_{W}^{\lambda}\quad\text{with $\theta(\lambda):=\lambda\log\lambda-\lambda+\frac{z}{2}\log\lambda$}.$
Introduce a differential operator $f(\lambda)$ in $\tau$ by
$f(\lambda):=\operatorname{rank}(\mathcal{Q})\theta(\lambda)-(\log\lambda)z\partial_{c_{1}(V)}$
where $\partial_{c_{1}(V)}$ denotes the directional derivative along $c_{1}(V)$ in the $\tau$-space. We consider the differential operator $f(\lambda+z\partial_{t})$ in both $\tau$ and $t$, defined via the Taylor expansion in $z\partial_{t}$, and define its regularization $f_{\rm reg}(\lambda,z\partial_{t})$ by subtracting terms involving $\log\lambda$:
$f_{\rm reg}(\lambda,z\partial_{t})$ $:=f(\lambda+z\partial_{t})-f(\lambda)-\operatorname{rank}(\mathcal{Q})(\log\lambda)z\partial_{t}$
$=\operatorname{rank}(\mathcal{Q})(\theta(\lambda+z\partial_{t})-\theta(\lambda)-(\log\lambda)z\partial_{t})-\log(1+\lambda^{-1}z\partial_{t})z\partial_{c_{1}(V)}.$
This belongs to $\mathbb{C}[z\partial_{t},z\partial_{c_{1}(V)}][\![\lambda^{-1}]\!]$ and satisfies $f_{\rm reg}(\lambda,z\partial_{t})|_{\lambda^{-1}=0}=0$.
The exact sequence (3.2) shows $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{\Delta}_{V}^{\lambda}=1$. Thus $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)$ is a $\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5. This is a miniversal slice, i.e. $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)|_{Q=\lambda^{-1}=0}=z+\tau+O(z^{-1})$. Therefore, we can apply Lemma 2.8 to this slice, and obtain a differential operator $\widetilde{F}(\lambda)\in\sum_{i=0}^{s}\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]z\partial_{\tau^{i}}$ such that $\widetilde{F}(\lambda)|_{Q=\lambda^{-1}=0}=0$ and
Lemma 2.7 and the QRR Theorem 2.5 imply that
is a $\mathbb{C}[\![Q,\tau,t,\lambda^{-1}]\!]$-valued point of the $(\mathcal{Q}(1),\tilde{e}_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$. We will calculate this element $\widetilde{I}^{\lambda}$. First observe the following identities:
(3.7) $\begin{split}&\widetilde{\Delta}^{\lambda}_{\mathcal{Q}(1)}=e^{(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+p)-\theta(\lambda)-p\log\lambda)+c_{1}(\mathcal{Q})\log(1+\frac{p}{\lambda}))/z}\widetilde{\Delta}^{\lambda+p}_{\mathcal{Q}},\\ &\widetilde{\Delta}_{\mathcal{Q}}^{\lambda+kz}/\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}=e^{-\left(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+kz)-\theta(\lambda))+c_{1}(\mathcal{Q})\log(1+\frac{kz}{\lambda})\right)/z}\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+\epsilon+cz),\\ \ &J_{V}^{\lambda+x}(\tau)\Bigr|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}=e^{-c_{1}(\mathcal{Q})\log(1+\frac{x}{\lambda})/z}e^{-\log(1+\frac{x}{\lambda})\partial_{c_{1}(V)}}\widetilde{J}_{V}^{\lambda+x}(\tau)\end{split}$
where $\prod_{\epsilon}$ is the product over the Chern roots of $\mathcal{Q}$. The first identity follows from $\Delta_{\mathcal{Q}(1)}^{\lambda}=\Delta_{\mathcal{Q}}^{\lambda+p}$, the second from (3.5), and the thrid from the Divisor Equation. We have
$\widetilde{I}^{\lambda}(\tau,t)$ $=(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}\sum_{k\geq 0}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}e^{f_{\rm reg}(\lambda,p+kz)/z}e^{\widetilde{F}(\lambda+p+kz)/z}J_{B}(\tau)$ by (3.3)
$=\sum_{k\geq 0}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}e^{f_{\rm reg}(\lambda,p+kz)/z}(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}\widetilde{\Delta}^{\lambda+p+kz}_{\mathcal{Q}}\widetilde{J}^{\lambda+p+kz}_{V}(\tau)$ by (3.6)
@@
$=\sum_{k\geq 0}e^{pt/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}\left(1+\frac{p+\epsilon+cz}{\lambda}\right)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)\Bigr|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ by (3.7).
By the change of variables $Q^{d}\to Q^{d}\lambda^{c_{1}(\mathcal{Q})\cdot d}=Q^{d}\lambda^{-c_{1}(V)\cdot d}$, $q^{k}\to q^{k}\lambda^{\operatorname{rank}(\mathcal{Q})k}$ for $\widetilde{I}^{\lambda}$, we obtain a point $I^{\lambda}$ of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone by Remark 2.6. It is given by
$I^{\lambda}(\tau,t)=\sum_{k\geq 0}e^{tp/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+p+\epsilon+cz)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau).$
The non-equivariant limit $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ exists and gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1. As discussed in §2.5, $I_{\mathbb{P}(V)}$ lies on the Givental cone of $\mathbb{P}(V)$ . The proof of Theorem 1.1 is now complete.
@@
The differential operator $f(\lambda)$ appearing in the proof above is chosen to satisfy the identity $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)$, which leads to
$\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}e^{\widetilde{F}(\lambda)/z}J_{B}(\tau).$
Conversely, if one chooses a differential operator $F(\lambda)=F(\lambda;\tau,z\partial_{\tau},z)$ in $\tau$ such that $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{F(\lambda)/z}J_{B}(\tau)$, then a straightforward calculation using (3.5) shows that the element $(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$ yields the point $I^{\lambda}(\tau,t)$ defined above. This point is expected to lie on the $(\mathcal{Q}(1),e_{\lambda})$-twisted cone of $B\times\mathbb{P}^{N-1}$, and the preceding construction makes this heuristic rigorous via the $\lambda^{-1}$-adic topology.
Remark 3.3.